# Quantum reservoir probing of quantum phase transitions

### v2

@@ Abstract
Quantum phase transitions are highly remarkable phenomena manifesting in quantum many-body systems. However, their precise identifications in equilibrium systems pose significant theoretical and experimental challenges. Thus far, dynamical detection protocols employing global quantum quenches have been proposed, wherein transitions are discerned from global nonequilibrium excitations. In this work, we demonstrate that quantum phase transitions can be detected through localized out-of-equilibrium excitations induced by local quantum quenches. While the resulting dynamics after the quench are influenced by both the local quench operation and the intrinsic dynamics of the quantum system, the effects of the former are exclusively extracted through the cutting-edge framework called quantum reservoir probing (QRP). Through the QRP, we find that the impacts of the local quenches vary across different quantum phases and are significantly suppressed by quantum fluctuations amplified near quantum critical points, thereby precisely delineating phase boundaries. We demonstrate that the QRP can detect quantum phase transitions in the paradigmatic integrable and nonintegrable quantum systems, and even topological quantum phase transitions, all within the identical framework employing single-site observables.
@@ INTRODUCTION
Quantum phase transitions emerge as quintessential manifestations of the interplay among quantum many-body phenomena [1, 2]. They are characterized by the universal description across microscopically different systems, which has been central to our understanding of quantum matters. Nevertheless, precisely detecting these transitions in equilibrium systems poses significant challenges both theoretically and experimentally. This has led recent investigations to focus on nonequilibrium dynamics, which is inherently related to quantum phase transitions through dynamical quantum critical phenomena. Indeed, nonequilibrium signatures of quantum phase transitions have been identified in a diverse range of systems, spanning integrable, nonintegrable, and topological systems [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. A prevalent strategy in these explorations involves a global quantum quench, wherein the entire system is initially configured in a specific state and subsequently undergoes time evolution under a Hamiltonian with suddenly changed parameters. Recent experimental advancements in isolating and manipulating quantum systems have facilitated such sophisticated control in various platforms, including trapped ions [18, 19, 20, 21], ultracold atoms [22, 23, 24, 25], Rydberg atoms [26, 27, 28, 29], and nitrogen-vacancy centers [30].
@@
A pertinent question then emerges regarding the feasibility of dynamically detecting equilibrium quantum phase transitions through local quantum quenches. Given that the local quench operation, in contrast to its global counterpart, selectively influences nontrivial excitations or quasiparticles in specific regions characteristic of the system [31, 32, 33, 34], the resulting dynamics may potentially unveil signatures of quantum phase transitions. However, the complexity lies in isolating and analyzing the response dynamics directly ascribed to the local quench operation. Since the dynamics is influenced by both the local quantum quench and the intrinsic dynamics of the quantum system itself, the latter obscures the former, thereby diminishing the potential benefit of local quenches in facilitating access to localized physics.
Figure 1: Conceptualization of the QRP. a Schematic of the QRC, which harnesses the quantum system as a computational resource. The quantum state of the reservoir reflects the provided input information, and the final output is obtained by linearly transforming a read-out vector constructed through multiple measurements on various degrees of freedom. b Schematic representation of the QRP, which investigates the quantum system from the perspective of information processing capabilities. The input is provided via the local quench, the output is calculated from a local degree of freedom, and the estimation performance of the input signifies the impact of the local quench on that degree of freedom. c Preparation and local manipulation to construct the initial state $|\Psi^{\mathrm{in}}_{k}\rangle$. The gray spin undergoes the local quench induced by the yellow needle, while the remaining spins are initialized in the designated state $|\Phi\rangle$, illustrated here as the all-up state. d Diagrammatic representation of the initial quantum state [Eq. (1)]. The central spin is locally manipulated contingent upon the input value $s_{k}$ for each case (three cases are depicted). This local quench induces varied dynamics in the local operator $O_{i}(t)$, which are reflective of the corresponding input values. e Dynamics of $\langle O_{i}(t)\rangle$ summarized across three cases depicted in (d). f Schematic illustration of the determination coefficient $R_{i}^{2}(t)$ derived from $\langle O_{i}(t)\rangle$. An almost unity $R^{2}_{i}(t)$ signifies a precise estimation of the input value, indicating the pronounced influence of the local quench on the observed $\langle O_{i}(t)\rangle$ [Eq. (4)]. Conversely, $R^{2}_{i}(t)$ approaches zero when $\langle O_{i}(t)\rangle$ exhibits similar values that do not reflect each input, as exemplified by intersections of $\langle O_{i}(t)\rangle$.
@@
In this paper, we reveal that signatures of quantum phase transitions manifest themselves in out-of-equilibrium local observables subsequent to local quantum quenches. To this end, we develop the methodology of “quantum reservoir probing” (QRP) [35], which selectively extracts the effects of the local quench from the dynamics influenced by multiple factors. Through the lens of the QRP, we observe that the local quantum quench exerts varied influences on a single-site operator depending on the inherent characteristics of quantum phases. Moreover, at the quantum critical points, these influences are notably suppressed due to the enhanced quantum fluctuations. These observations highlight the capability of our QRP to effectively detect quantum phase transitions, as exemplified in both integrable and nonintegrable models herein. Furthermore, notwithstanding the use of a single-site operator, we demonstrate that the QRP can identify topological quantum phase transitions, which are typically distinguished by the absence of local order parameters beyond the Landau-Gintzburg-Wilson paradigm. The broad applicability and design flexibility of our QRP establish it as a versatile tool for probing quantum phenomena, opening the door to further understanding of quantum many-body physics.
@@ RESULTS
The QRP is an inverse extension of the QRC, wherein the quantum system employed as the reservoir part is investigated by linking physical phenomena with information processing (Fig. 1b). As introduced below, the input is provided via the local quench, and the output is calculated from a local observable; the calculation objective is to estimate the original input value. Given that information is provided solely through the local quench, the estimation succeeds when the local quench exerts an influence on the observed degree of freedom, whereas it fails when the operator is entirely unaffected by the local quench. By systematically scanning local observation operators, the QRP selectively probes the impacts of the local quench on the quantum system through the input estimation performance.
@@
To elaborate, the central feature underlying the QRP is the manipulation of a local degree of freedom, parametrized by a random input value $s_{k}$. Our focus is on a local operator $O$ defined per site $i$ and time $t$, denoted as $O_{i}(t)$, with particular attention to the variations induced by the input variable $s_{k}$. In the statistical analysis across various instances with different inputs subscripted by $k$, the expectation value $\langle O_{i}(t)\rangle_{k}$ may display diverse characteristics that closely reflect the corresponding input values. In such cases, the value of $s_{k}$ can be precisely deduced from $\langle O_{i}(t)\rangle_{k}$ through straightforward calculations. Hence, the successful estimation of the original input value $s_{k}$ provides evidence that the local manipulation exerts an influence on the operator $O_{i}(t)$ at that particular spatiotemporal coordinate. In contrast, the operator $O_{j}(t^{\prime})$ at a different spatiotemporal point may demonstrate consistent behavior across all instances, not accurately mirroring the local quench operation. In this latter case, the operator is predominantly governed by the inherent characteristics of the quantum system independent of the local quench. Therefore, the dependency of the operator $O_{i}(t)$ on the input value $s_{k}$ explores how the local quantum quench influences the operator, while isolating it from other intrinsic phenomena. Remarkably, this linkage is highly sensitive to the Hamiltonian governing the quantum system. The operator’s dependency on the local quench operation thus serves as a witness to the properties of the quantum system, shedding light on the underlying quantum processes at play.
@@
Here, we explain the specific procedures of the QRP on the one-dimensional (1D) quantum systems. The protocol consists of four steps: (i) the preparation of the initial state (Fig. 1c), (ii) the local quench on the central spin, parameterized by the input value $s_{k}$ (Figs. 1c and 1d), (iii) the observation of the dynamics of a local operator $\langle O_{i}(t)\rangle$ under the Hamiltonian $\mathcal{H}$ (Figs. 1d and 1e), and (iv) the statistical evaluation of the performance in estimating the input value $s_{k}$ from the observed dynamics, quantified via the determination coefficient $R_{i}^{2}(t)$ (Fig. 1f). We provide a comprehensive description of each step in the following.
@@
First, the system is prepared in a specified initial state in step (i); for simplicity, a product state is assumed here (e.g., the all-up state in Fig. 1c). Subsequently, in step (ii), the central spin undergoes local manipulation based on the input value $s_{k}$. The resultant state is expressed as
@@
where $|\psi(s_{k})\rangle$ represents the state of the locally manipulated spin, while $|\Phi\rangle$ denotes the prepared initial state for the remaining subsystem (Fig. 1d). The input value $s_{k}$ is randomly sampled from a uniform distribution, specifically $s_{k}\in[0,1]$. For the spin-$1/2$ case, we consider the initial state as
@@
where |±=(±$|\pm\rangle=\left(\mid\uparrow\rangle\pm\mid\downarrow\rangle\right)/\sqrt{2}$. This state can be prepared by applying a single-gate pulse or a local magnetic field $\bm{h}\propto(2s_{k}-1,0,-2\sqrt{s_{k}(1-s_{k})})$.
@@
where $\mathrm{cov}$ and $\sigma^{2}$ represent covariance and variance, respectively; see Methods section for more details. The value of $R^{2}_{i}(t)$ ranges between 00 and $1$, with an approach towards $1$ indicating that the outputs $\{y_{{i,k}}(t)\}$ closely reproduce the inputs $\{s_{k}\}$ for all testing cases.
Figures 1e and 1f visually elucidate the relationship between the dynamics of $\langle O_{i}(t)\rangle_{k}$ and the corresponding $R^{2}_{i}(t)$. A high $R^{2}_{i}(t)$ value suggests that $\langle O_{i}(t)\rangle_{k}$ varies systematically in response to the input value $s_{k}$. In contrast, $R^{2}_{i}(t)$ close to 00 indicates that $\langle O_{i}(t)\rangle_{k}$ is largely independent of $s_{k}$, such as at the crossing points resulting from intrinsic oscillations in the quantum system. Therefore, $R^{2}_{i}(t)$ serves as an indicator of the extent to which $O_{i}(t)$ is influenced by the local quench operation. Under different parameters of the Hamiltonian $\mathcal{H}$, the local quench operation exerts varying influences on the resultant dynamics, hence the behavior of $R^{2}_{i}(t)$ is instrumental in examining the unique properties of each quantum phase. It is worth emphasizing that restricting the transformation generating $y_{{i,k}}(t)$ from $\langle O_{i}(t)\rangle_{k}$ to a linear form is crucial. Any additional nonlinearity could lead to either an overestimation or underestimation of the effects of the local quench, while simultaneously impeding interpretability as the output is influenced by the extraneous characteristics of the transformation itself.
Figure 2: Identifying the quantum phase transition in the transverse-field Ising model via the QRP. a Spatiotemporal representation of the performance for the estimation of the input value. The color represents the value of $R^{2}_{i}(t)$ derived from $\langle\sigma^{x}_{i}(t)\rangle$. The upper band represents the schematic phase diagram regarding $\mathcal{H}(g)$, where the quantum critical point separating the ferromagnetically ordered phase and the quantum disordered phase exists at ${g_{c}}=1$. The system size is $N=49$, and the time evolution is calculated using a bond dimension $\chi=256$; the result is quantitatively the same for $\chi=128$. b The early-time spin dynamics at the nearest-neighbor site from the central one under three different input values. c The estimation performance $R^{2}_{i}(t)$ utilizing $\langle\sigma^{x}_{i}(t)\rangle$ at the same site as (b). d The average $\overline{R^{2}}$ over spatiotemporal indices $(i,t)\in\Lambda\times\{0<t\leq 10\}$, with $\Lambda$ encompassing the central $11$ sites. The blue and pink lines represent the system sizes $N=49$ and $N=37$, respectively. The quantum critical point is denoted by the gray dashed line.
@@
The system is initialized in the all-up state. Following the local quantum quench, the resultant state $|\Psi^{\mathrm{in}}_{k}\rangle$ in Eq. (1) is explicitly expressed as:
$|\Psi^{\mathrm{in}}_{k}\rangle=\mid\uparrow\uparrow\uparrow\dots\rangle\otimes|\psi(s_{k})\rangle\otimes\mid\uparrow\uparrow\uparrow\dots\rangle.$ (6)
@@
Subsequently, we monitor the dynamics of the local spin $\langle\sigma^{x}_{i}(t)\rangle$ at each site $i$ under the Hamiltonian $\mathcal{H}=\mathcal{H}(g)$. Repeating the procedure for many instances with varying $s_{k}$, the estimation performance $R^{2}_{i}(t)$ in Eq. (4) is methodically evaluated.
@@
Figure 2 encapsulates the results of the QRP under various magnetic fields $g$. In Fig. 2a, we show the spatiotemporal representation of the estimation performance $R^{2}_{i}(t)$, which reveals a discernible light-cone structure emanating from the quenched central site. The broad distribution of nonzero $R^{2}_{i}(t)$ suggests that the local spin operator $\langle\sigma_{i}^{x}(t)\rangle$ across diverse spatial locations and temporal moments adequately reflect the impact of the local quantum quench. In the quantum disordered phase, where the ground state exhibits spins aligned along the $x$ direction, the effects of the local quench along the $x$ axis in spin space propagate akin to a spin wave, characterized by pronounced wavefronts of $R^{2}_{i}(t)$. In contrast, in the ferromagnetically ordered phase, the local quenching effects are preserved predominantly within the quasiparticles, manifesting as alterations in phase and amplitude. This influence is evidenced by the ripple-like patterns of the nonzero $R^{2}_{i}(t)$. An exception occurs near the quantum critical point at $g_{c}=1$, characterized by maximized fluctuations that lead to an effective breakdown of the quasiparticle picture. Here, the effects of the local quench are passed on the intrinsic fluctuations of the system, though such influences rapidly diminish as indicated by the noticeably small $R^{2}_{i}(t)$ after the passage of the wavefront. This observation highlights an anomalous suppression of the propagation of local quenching effects at the quantum critical point, which is distinct from other non-critical phases.
@@
In Fig. 2b, the early-time dynamics of $\langle\sigma^{x}_{i}(t)\rangle$ at the nearest-neighbor spin from the central one is depicted for three distinct input values $s_{k}$ at different magnetic fields $g$. At $g=0.5$ and $1.5$, the quasiparticles induce periodic spin oscillations, the phase and amplitude of which are dependent on the input $s_{k}$. In contrast, at $g=1.0$, the dynamics of $\langle\sigma^{x}_{i}(t)\rangle$ exhibits significant variation when the wavefronts pass at approximately $t\simeq 1$, thereafter stabilizing into a uniform pattern that does not reflect the influence of quench operation. Figure 2c displays the corresponding estimation performance at the same site (a vertical cross-section of the color map in Fig. 2a). Over broad temporal spans at $g=0.5$ and $1.5$, $R^{2}_{i}(t)$ becomes close to $1$, which statistically underscores the input dependency of the phase and amplitude of the spin oscillations. Notably, their frequency is governed not by the local quench on the initial state, but rather by the intrinsic characteristics of the quasiparticles themselves. This is evidenced by $R^{2}_{i}(t)$ periodically approaching 00 (Fig. 2c) at points where oscillations of $\langle\sigma_{i}^{x}(t)\rangle$ intersect for different inputs (Fig. 2b). In contrast, at $g=1.0$, while $R^{2}_{i}(t)$ nears $1$ around $t\simeq 1$ where the wavefronts induce variations in the spin dynamics, the similarity of $\langle\sigma_{i}^{x}(t)\rangle$ in other temporal regions leads to $R^{2}_{i}(t)\simeq 0$, signifying the reduced influence of the local quench by the intrinsic fluctuations enhanced near the quantum critical point.
@@
As described above, the effects of the local quantum quench on the resultant quantum dynamics of local operators manifest qualitatively different characteristics across various phases (Fig. 2a). The distribution of $R^{2}_{i}(t)$ thus serves as a marker for identifying the quantum phase transition. To enhance the clarity of the results, in Fig. 2d, we display the mean value of $R^{2}_{i}(t)$ over a specific subset of spatiotemporal indices $\{(i,t)\}$, denoted as $\overline{R^{2}}$. This analysis focus on $(i,t)$ within the central eleven sites during the time interval $0<t\leq 10$ to minimize the effects from the edges of the system; selecting different subsets does not qualitatively alter the results (see the Supplemental Information). We note that $\overline{R^{2}}$ represents a simple average of local quantities, and therefore, does not include nonlocal quantum effects. As $g$ approaches $g_{c}=1$ from the ferromagnetically ordered state, the fluctuations progressively intensify and interfere with the estimation of the original input, resulting in a decrease in $\overline{R^{2}}$. At the quantum critical point, $\overline{R^{2}}$ reaches its minimum, which corresponds to the emergence of a dark region between the wavefronts observed in Fig. 2a for $g=1.0$. Upon further augmentation of $g$ into the quantum disordered phase, the fluctuations weaken again and the effects of the quench operation become dominant, leading to an increase in $\overline{R^{2}}$. The pronounced dip in Fig. 2d is therefore a definitive indicator of the quantum phase transition occurring at $g_{c}$. Consequently, by focusing on the effects of the local quench, our QRP successfully detects the quantum phase transition only by measurements of single-site spin dynamics.

### v3

@@ Abstract
Quantum phase transitions are highly remarkable phenomena manifesting in quantum many-body systems. However, their precise identifications in equilibrium systems pose significant theoretical and experimental challenges. Thus far, dynamical detection protocols employing global quantum quenches have been proposed, wherein transitions are discerned via global nonequilibrium excitations. In this work, we demonstrate that quantum phase transitions can be detected through localized out-of-equilibrium excitations induced by local quantum quenches. While the resulting dynamics after the quench is influenced by both the local quench operation and the intrinsic dynamics of the quantum system, the effects of the former are exclusively extracted using the cutting-edge framework called quantum reservoir probing (QRP). Through the QRP, we find that the impacts of the local quenches vary across different quantum phases and are significantly suppressed by quantum fluctuations amplified near quantum critical points; consequently, phase boundaries are precisely delineated. We demonstrate that the QRP can detect quantum phase transitions in the paradigmatic integrable and nonintegrable quantum spin systems, and even topological quantum phase transitions, all within the identical framework employing local quantum quenches and single-site observables.
@@ INTRODUCTION
Quantum phase transitions emerge as quintessential manifestations of the interplay among quantum many-body phenomena [1, 2]. They are characterized by the universal description across microscopically different systems, which has been central to our understanding of quantum matters. Nevertheless, precisely detecting these transitions in equilibrium systems poses significant challenges both theoretically and experimentally. This has led recent investigations to focus on nonequilibrium dynamics, which is inherently related to quantum phase transitions through dynamical quantum critical phenomena. Indeed, nonequilibrium signatures of quantum phase transitions have been identified in a diverse range of systems, spanning integrable, nonintegrable, and topological systems [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. A prevalent strategy in these explorations involves a global quantum quench, wherein the entire system is initially configured in a designated state and subsequently undergoes time evolution under a specific Hamiltonian. Recent experimental advancements in isolating and manipulating quantum systems have facilitated such sophisticated control in various platforms, including trapped ions [18, 19, 20, 21], ultracold atoms [22, 23, 24, 25], Rydberg atoms [26, 27, 28, 29], and nitrogen-vacancy centers [30].
@@
A pertinent question then emerges regarding the feasibility of dynamically detecting equilibrium quantum phase transitions through local quantum quenches. Given that the local quench operation, in contrast to its global counterpart, selectively influences nontrivial excitations or quasiparticles in specific regions characteristic of each quantum phase [31, 32, 33, 34], the resulting dynamics may potentially unveil signatures of quantum phase transitions. However, the complexity lies in isolating and analyzing the response dynamics directly ascribed to the local quench operation. Since the dynamics is influenced by both the local quantum quench and the intrinsic dynamics of the quantum system itself, the latter obscures the former, thereby diminishing the potential benefit of local quenches in facilitating access to localized physics.
@@
In this paper, we reveal that signatures of quantum phase transitions manifest themselves in out-of-equilibrium local observables subsequent to local quantum quenches. To this end, we develop the methodology of “quantum reservoir probing” (QRP) [35], which selectively extracts the effects of the local quantum quench from the dynamics influenced by multiple factors. Through the lens of the QRP, we observe that the local quench exerts varied influences on a single-site operator, depending on the inherent characteristics of quantum phases. Moreover, at the quantum critical points, these influences are notably suppressed due to the enhanced quantum fluctuations. Capitalizing on these phenomena, the QRP demonstrates the robust capability to detect quantum phase transitions, as exemplified in both integrable and nonintegrable models herein. Furthermore, notwithstanding the use of local quenches and single-site operators, we demonstrate that the QRP can identify topological quantum phase transitions, which are typically distinguished by the absence of local order parameters beyond the Landau-Ginzburg-Wilson paradigm. The broad applicability and design flexibility of our QRP establish it as a versatile tool for probing a variety of quantum phenomena, opening the door to further understanding of quantum many-body physics.
@@ RESULTS
Figure 1: Conceptualization of the QRP. a Schematic of the QRC, which harnesses the quantum system as a computational resource. The quantum state of the reservoir reflects the provided input information, and the final output is obtained by linearly transforming a read-out vector constructed through multiple measurements on various degrees of freedom. b Schematic representation of the QRP, which investigates the quantum system from the perspective of information processing capabilities. The input is provided via the local quench, the output is calculated from a local degree of freedom, and the estimation performance of the input signifies the impact of the local quench on that degree of freedom. c Preparation and local manipulation to construct the initial state $|\Psi^{\mathrm{in}}_{k}\rangle$. The gray spin undergoes the local quantum quench induced by the yellow needle, while the remaining spins are initialized in the designated state $|\Phi\rangle$, illustrated here as the all-up state. d Diagrammatic representation of the initial quantum state [Eq. (1)]. The central spin is locally manipulated contingent upon the input value $s_{k}$ for each case (three instances are depicted). This local quench induces varied dynamics in the local operator $O_{i}(t)$, which is reflective of the corresponding input values. e Dynamics of $\langle O_{i}(t)\rangle$ summarized across three cases depicted in (d). f Schematic illustration of the determination coefficient $R_{i}^{2}(t)$ derived from $\langle O_{i}(t)\rangle$. An almost unity $R^{2}_{i}(t)$ signifies a precise estimation of the input value, indicating the pronounced influence of the local quench on the observed $\langle O_{i}(t)\rangle$ [Eq. (4)]. Conversely, $R^{2}_{i}(t)$ approaches zero when $\langle O_{i}(t)\rangle$ exhibits similar values that do not reflect each input, as exemplified by intersections of $\langle O_{i}(t)\rangle$.
@@
The QRP is an inverse extension of the QRC, wherein the quantum system employed as the reservoir part is investigated by linking physical phenomena with information processing (Fig. 1b). As introduced below, the input is provided via the local quantum quench, and the output is computed from a local observable; the objective of the calculation is to estimate the original input value. Given that information is provided solely through this local quench, the estimation succeeds when the local quench exerts an influence on the observed degree of freedom. By systematically scanning local observation operators, the QRP selectively probes the impacts of the local quench on the quantum system through the input estimation performance.
@@
To elaborate, the central feature underlying the QRP is the manipulation of a local degree of freedom, parametrized by a random input value $s_{k}$. Our focus is on a local operator $O$ defined at spatial position site $i$ and time $t$, denoted as $O_{i}(t)$, with particular attention to the variations induced by the input variable $s_{k}$. In the statistical analysis across various instances with different inputs subscripted by $k$, the expectation value $\langle O_{i}(t)\rangle_{k}$ may display diverse characteristics that closely reflect the corresponding input values. In such cases, the value of $s_{k}$ can be precisely deduced from $\langle O_{i}(t)\rangle_{k}$ through straightforward calculations. Hence, the successful estimation of the original input value $s_{k}$ provides evidence that the local manipulation exerts an influence on the operator $O_{i}(t)$ at that particular spatiotemporal coordinate. In contrast, when the operator $O_{j}(t^{\prime})$ at a different spatiotemporal point demonstrates consistent behavior across all instances, it is predominantly governed by the inherent characteristics of the quantum system independent of the local quench. Therefore, the dependency of the operator $O_{i}(t)$ on the input value $s_{k}$ explores how the local quantum quench influences the operator, while isolating it from other intrinsic phenomena. Remarkably, this linkage is highly sensitive to the Hamiltonian governing the quantum system. The operator’s dependency on the local quench operation thus serves as a witness to the properties of the quantum system, shedding light on the underlying quantum processes at play.
@@
Here, we explain the specific procedures of the QRP on the one-dimensional (1D) quantum systems. The protocol consists of four steps: (i) the preparation of the initial state (Fig. 1c), (ii) the local quench on the central spin, parameterized by the input value $s_{k}$ (Figs. 1c and 1d), (iii) the observation of the dynamics of a local operator $\langle O_{i}(t)\rangle$ under the Hamiltonian $\mathcal{H}$ (Figs. 1d and 1e), and (iv) the statistical evaluation of the performance in estimating the input value $s_{k}$ from the observed dynamics, quantified by the determination coefficient $R_{i}^{2}(t)$ (Fig. 1f). We provide a comprehensive description of each step in the following.
@@
First, the system is prepared in a designated initial state in step (i); for simplicity, a product state is assumed here (e.g., the all-up state in Fig. 1c). Subsequently, in step (ii), a specific spin undergoes a local quench manipulation based on the input value $s_{k}$. The resultant state is expressed as
@@
where $|\psi(s_{k})\rangle$ represents the state of the locally manipulated spin, while $|\Phi\rangle$ denotes the prepared initial state for the remaining subsystem (Fig. 1d). The input value $s_{k}$ is randomly sampled from a uniform distribution, specifically $s_{k}\in[0,1]$. For instance, we consider the initial state as
@@
where |±=(±$|\pm\rangle=\left(\mid\uparrow\rangle\pm\mid\downarrow\rangle\right)/\sqrt{2}$. This state can be prepared by applying a single-gate pulse or a local magnetic field $\mathbf{h}\propto(2s_{k}-1,0,-2\sqrt{s_{k}(1-s_{k})})$.
@@
where $\mathrm{cov}$ and $\sigma^{2}$ represent covariance and variance, respectively; see Methods section for more details. The value of $R^{2}_{i}(t)$ ranges between 00 and $1$, with an approach towards $1$ indicating that the outputs $\{y_{{i,k}}(t)\}$ closely reproduce the inputs $\{s_{k}\}$ for all testing instances.
Figure 2: Identifying the quantum phase transition in the transverse-field Ising model via the QRP. a Spatiotemporal representation of the performance for the estimation of the input value. The color represents the value of $R^{2}_{i}(t)$ derived from $\langle\sigma^{x}_{i}(t)\rangle$. The upper band represents the schematic phase diagram regarding $\mathcal{H}(g)$, where the quantum critical point separating the ferromagnetically ordered phase and the quantum disordered phase exists at ${g_{c}}=1$. The system size is $N=49$, and the time evolution is calculated using a bond dimension $\chi=256$; the result is quantitatively the same for $\chi=128$. b The early-time spin dynamics at the nearest-neighbor site from the central one under three different input values. c The estimation performance $R^{2}_{i}(t)$ utilizing $\langle\sigma^{x}_{i}(t)\rangle$ at the same site as (b). d The average $\overline{R^{2}}$ over spatiotemporal indices $(i,t)\in\Lambda\times\{0<t\leq 7\}$, with $\Lambda$ encompassing the central nine sites. The blue and pink lines represent the results for system sizes $N=49$ and $N=37$, respectively, exhibiting analogous behavior. The quantum critical point is denoted by the gray dashed line.
Figures 1e and 1f visually elucidate the relationship between the dynamics of $\langle O_{i}(t)\rangle_{k}$ and the corresponding $R^{2}_{i}(t)$. A high $R^{2}_{i}(t)$ value suggests that $\langle O_{i}(t)\rangle_{k}$ varies systematically in response to the input value $s_{k}$. In contrast, $R^{2}_{i}(t)$ close to 00 indicates that $\langle O_{i}(t)\rangle_{k}$ is largely independent of $s_{k}$, such as at the crossing points resulting from intrinsic oscillations in the quantum system. Therefore, $R^{2}_{i}(t)$ serves as an indicator of the extent to which $O_{i}(t)$ is influenced by the local quench operation. Under different parameters of the Hamiltonian $\mathcal{H}$, the local quench operation exerts varying influences on the resultant dynamics, hence the behavior of $R^{2}_{i}(t)$ is instrumental in examining the unique properties of each quantum phase. It is worth noting that, although the input in Eq. (2) includes a nonlinear contribution from $s_{k}$, our analysis concentrates on the spatiotemporal dynamics of the linear component in accordance with the QRC framework. Any additional complexity beyond linear regression could impede interpretability, as the output is influenced by the extraneous characteristics of the transformation itself.
@@
In step (i), the system is initialized in the all-up state, and after the local quantum quench at the central spin in step (ii), the resultant state $|\Psi^{\mathrm{in}}_{k}\rangle$ in Eq. (1) is explicitly expressed as
$|\Psi^{\mathrm{in}}_{k}\rangle=\mid\uparrow\uparrow\uparrow\dots\rangle\otimes|\psi(s_{k})\rangle\otimes\mid\uparrow\uparrow\uparrow\dots\rangle,$ (6)
@@
using $|\psi(s_{k})\rangle$ defined in Eq. (2). In step (iii), we monitor the dynamics of the local spin $\langle\sigma^{x}_{i}(t)\rangle$ at each site $i$ under the Hamiltonian $\mathcal{H}=\mathcal{H}(g)$. Repeating the procedure for many instances with varying $s_{k}$, the estimation performance $R^{2}_{i}(t)$ in Eq. (4) is methodically evaluated in step (iv).
@@
Figure 2 encapsulates the results of the QRP under various magnetic fields $g$. In Fig. 2a, we show the spatiotemporal representation of the estimation performance $R^{2}_{i}(t)$, which reveals a discernible light-cone structure emanating from the quenched central site. The broad distribution of nonzero $R^{2}_{i}(t)$ suggests that the local spin operator $\langle\sigma_{i}^{x}(t)\rangle$ across diverse spatial locations and temporal moments adequately reflect the impact of the local quantum quench. In the quantum disordered phase, where the ground state exhibits spins aligned along the $x$ direction, the effects of the local quench along the $x$ axis in spin space propagate akin to a spin wave, characterized by pronounced wavefronts of $R^{2}_{i}(t)$. In contrast, in the ferromagnetically ordered phase, the local quenching effects are preserved predominantly within the quasiparticles, manifesting as alterations in phase and amplitude. This influence is evidenced by the ripple-like patterns of the nonzero $R^{2}_{i}(t)$. An exception occurs near the quantum critical point at $g_{c}=1$, characterized by maximized fluctuations that lead to an effective breakdown of the quasiparticle picture. Here, the effects of the local quench are passed on the intrinsic fluctuations of the system, though such influences rapidly diminish as indicated by the noticeably small $R^{2}_{i}(t)$ after the passage of the wavefront. This observation highlights an anomalous suppression of the local quenching effects at the quantum critical point, which is distinct from those in other non-critical phases.
@@
In Fig. 2b, the early-time dynamics of $\langle\sigma^{x}_{i}(t)\rangle$ at the nearest-neighbor spin from the central one is depicted for three distinct input values $s_{k}$ at different magnetic fields $g$. At $g=0.5$ and $1.5$, periodic spin oscillations are induced, with their phase and amplitude dependent on the input $s_{k}$. In contrast, at $g=1.0$, the dynamics of $\langle\sigma^{x}_{i}(t)\rangle$ exhibits significant variation when the wavefronts pass at approximately $t\simeq 1$, thereafter stabilizing into a featureless pattern with diminished influence from the quench operation. Figure 2c displays the corresponding estimation performance at the same site (a vertical cross-section of the color map in Fig. 2a). Over broad temporal spans at $g=0.5$ and $1.5$, $R^{2}_{i}(t)$ becomes close to $1$, which statistically underscores the input dependency of the phase and amplitude of the spin oscillations. In contrast, their frequency is governed not by the local quench on the initial state, but rather by the intrinsic characteristics of the quasiparticles themselves. This is evidenced by $R^{2}_{i}(t)$ periodically approaching zero (Fig. 2c) at points where oscillations of $\langle\sigma_{i}^{x}(t)\rangle$ intersect for different inputs (Fig. 2b). Notably, at $g=1.0$, while $R^{2}_{i}(t)$ nears $1$ around $t\simeq 1$ where the wavefronts induce variations in the spin dynamics, $R^{2}_{i}(t)\simeq 0$ in other temporal regions signify the reduced influence of the local quench due to the intrinsic fluctuations enhanced near the quantum critical point.
@@
As described above, the effects of the local quantum quench on the resultant dynamics of local operators manifest qualitatively different characteristics across various phases (Fig. 2a). The distribution of $R^{2}_{i}(t)$ thus serves as a marker for identifying the quantum phase transition. To enhance the clarity of the results, in Fig. 2d, we display the mean value of $R^{2}_{i}(t)$ over a specific subset of spatiotemporal indices $\{(i,t)\}$, denoted as $\overline{R^{2}}$. This analysis focuses on $(i,t)$ within the central nine sites during the time interval $0<t\leq 7$; selecting different subsets does not qualitatively alter the results (see the Supplementary Information). We note that $\overline{R^{2}}$ represents a simple average of local quantities, and therefore, does not include nonlocal quantum effects. As $g$ approaches $g_{c}=1$ from the ferromagnetically ordered state, the fluctuations progressively intensify and interfere with the estimation of the original input, resulting in a decrease in $\overline{R^{2}}$. At the quantum critical point, $\overline{R^{2}}$ reaches its minimum, which corresponds to the emergence of a dark region between the wavefronts observed in Fig. 2a for $g=1.0$. Upon further augmentation of $g$ into the quantum disordered phase, the fluctuations weaken again and the effects of the quench operation become dominant, leading to an increase in $\overline{R^{2}}$. The pronounced dip in Fig. 2d is therefore a definitive indicator of the quantum phase transition occurring at $g=g_{c}$. Consequently, by focusing on the effects of the local quantum quench, our QRP successfully detects the quantum phase transition only by measurements of single-site spin dynamics.
Figure 3: Detection of the quantum phase transition in the ANNNI model using the QRP. a Spatiotemporal representation of the estimation performance when employing $\langle\sigma^{x}_{i}(t)\rangle$ in the nonintegrable ANNNI chain characterized by the interaction parameter $\kappa=0.5$. The system size is $N=49$, and the time evolution is calculated with a bond dimension of $\chi=256$. b The average $\overline{R^{2}}$ over a subset $(i,t)\in\Lambda\times\{0<t\leq 5\}$, where $\Lambda$ contains the central seven sites. The upper band represents the schematic phase diagram, where the ferromagnetically ordered and quantum disordered phases are separated by the quantum critical point at approximately $g_{c}\simeq 1.7$.